For f(x) = |x-5| and g(x) = 2x + 1, find
a. (f °g)(1) b. (g °g)(1)

Answer by stanbon(75887)

For f(x) = |x-5| and g(x) = 2x + 1, find
a. (f °g)(1) = f[g(1)] = f[2+1] = f(3) = |3-5| = 2
-------------------------
b. (g °g)(1) = g[g(1)] = g[3] = 2*3+1 = 7
--------------
Cheers,
Stan H.
----------